OpenStudy (anonymous):

find the function value: f(3)-9=5n squared + 4n

OpenStudy (anonymous):

The idea here, is to replace all n's by 3. Let me give you a similar example: f(4) = 3m^2 +2n, will be: f(4) = 3(4)^2 + 2(4) = 3(16) + 8 = 48 +8 = 56.